A boat travels 12 km. upstream from town A to town B in 6 hours. The boat
then travels downstream from town B to town C, which is 8 km. further downstream
from A in 5 hours. What is the speed of the boat in still water?

Found 2 solutions by nyc_function, mananth:
Answer by nyc_function(2741)

Going upstreams slows the boat; going downstream will speed up the boat.
Boat Upstream:
rate = x - 12
time = 6
Boat Downstream:
rate = x + 8
time = 5
Distances are equal.
Let x = speed of the boat in still water

6(x - 12) = 5(x + 8)
6x - 72 = 5x + 40
6x - 5x = 72 + 40
x = 112 km per hour.

Answer by mananth(16946)

let the speed of boat in still water be x km/h
speed of current = y km/h
..
distance =12 km
speed = x-y
time = 12/(x-y)
..
downstream
distance = 20km
speed = x+y
time = 20/x+y
..
12/(x-y)= 6
20/(x+y)= 5
6x-6y=12
x-y=2..................1
..
5x+5y= 20
x+y=4..................2
add 1 & 2
2x=6
divide by 2
x=3 speed of boat in still water
x-y=2
3-y=2
y=1 current speed